# VaR at Its Extremes: Impossibilities and Conditions for One-Sided Random Variables

### v2

@@ 3 VaR super-additivity
    $\phi_{i}^{\prime}(x_{i})$ $=\alpha_{i}\left(\log\!\left(\frac{\displaystyle x_{i}}{\displaystyle 1+x_{i}}\right)+\frac{\displaystyle 1}{\displaystyle 1+x_{i}}\right)$
    $=\alpha_{i}\left(-\sum_{k=1}^{\infty}\frac{1}{k}\left(\frac{1}{1+x_{i}}\right)^{k}+\frac{\displaystyle 1}{\displaystyle 1+x_{i}}\right)$
    $\leq\alpha_{i}\left(-\frac{1}{1+x_{i}}+\frac{\displaystyle 1}{\displaystyle 1+x_{i}}\right)=0.$
@@
    Log-hazard distribution. If
    $F_{X_{i}}(x_{i})=\exp\!\left(-\frac{\displaystyle\log(1+x_{i})^{\alpha_{i}}}{\displaystyle x_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i}\in(-\infty,1),$
@@
    $F_{X_{i}}(x_{i})=\frac{\displaystyle 1}{\displaystyle 2}+\frac{\displaystyle 1}{\displaystyle\pi}\arctan(\alpha_{i}\log(x_{i})),\qquad x_{i}\geq 0,\ \alpha_{i}>0.$
@@
    $\phi_{i}(x_{i})=x_{i}\log\!\left(\frac{\displaystyle 1}{\displaystyle 2}+\frac{\displaystyle 1}{\displaystyle\pi}\arctan(\alpha_{i}\log(x_{i}))\right).$
@@
    $\phi_{i}^{\prime}(x_{i})=\log F_{X_{i}}(x_{i})+\frac{\displaystyle\alpha_{i}}{\displaystyle\pi(1+(\alpha_{i}\log(x_{i}))^{2})\,F_{X_{i}}(x_{i})}.$
@@
    $\phi_{i}^{\prime}(x_{i})=\log F_{X_{i}}(x_{i})+\frac{\displaystyle\alpha_{i}\cos^{2}\theta}{\displaystyle\pi F_{X_{i}}(x_{i})}=\log F_{X_{i}}(x_{i})+\frac{\displaystyle\alpha_{i}\sin^{2}(\pi F_{X_{i}}(x_{i}))}{\displaystyle\pi F_{X_{i}}(x_{i})}.$
@@
    $\psi_{i}(u)=\log(u)+\frac{\displaystyle\alpha_{i}\sin^{2}(\pi u)}{\displaystyle\pi u},\qquad u\in[0,1].$
@@
    $0<\alpha_{i}\leq\inf_{u\in[0,1]}\frac{\displaystyle-\pi u\log(u)}{\displaystyle\sin^{2}(\pi u)}\approx 1.0568.$
@@
    $F_{X_{i}}(x_{i})=\frac{1}{\Gamma(\alpha_{i})}\Gamma\!\left(\alpha_{i},\frac{\displaystyle\theta_{i}}{\displaystyle x_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i},\theta_{i}>0,$
@@
    $\phi_{i}(x_{i})=x_{i}\log\!\left(\frac{1}{\Gamma(\alpha_{i})}\Gamma(\alpha_{i},\tfrac{\theta_{i}}{x_{i}})\right).$
@@
    $\phi_{i}^{\prime}(x_{i})=\log F_{X_{i}}(x_{i})+\frac{\displaystyle\left(\frac{\displaystyle\theta_{i}}{\displaystyle x_{i}}\right)^{\alpha_{i}}\exp(-\theta_{i}/x_{i})}{\displaystyle\Gamma(\alpha_{i})\,F_{X_{i}}(x_{i})}.$
@@
    $\psi_{i}(t)=\log\!\left(\frac{1}{\Gamma(\alpha_{i})}\Gamma(\alpha_{i},t)\right)+\frac{\displaystyle t^{\alpha_{i}}\exp(-t)}{\displaystyle\Gamma(\alpha_{i},t)}.$
@@
    a Gautschi-type lower bound [Gautschi1959]. Thus $\psi_{i}$ is non-increasing with $\psi_{i}(0)=0$, proving $\psi_{i}(t)\leq 0$ on $0,\infty)$.
@@
    $x_{i}h_{X_{i}}(x_{i})\leq\int_{x_{i}}^{\infty}h_{X_{i}}(w)\mathrm{d}w,\quad\forall x_{i}\in[0,\infty).$
    Where $h_{X_{i}}(x_{i})=\frac{\displaystyle f_{X_{i}}(x_{i})}{\displaystyle F_{X_{i}}(x_{i})}$ is the reverse hazard rate function [Block1998] of the random variable $X_{i}$.
@@
    $G_{i}(\lambda x_{i})\leq\frac{\displaystyle 1}{\displaystyle\lambda}G(x_{i}),\ \quad\forall\lambda\in[1,\infty).$
@@
    $\phi_{i}^{{}^{\prime}}(x_{i})$ $=\log F_{X_{i}}(x_{i})+\frac{\displaystyle x_{i}f_{X_{i}}(x_{i})}{\displaystyle F_{X_{i}}(x_{i})},$
    $=-\int_{x_{i}}^{\infty}\frac{\displaystyle f_{X_{i}}(w)}{\displaystyle F_{X_{i}}(w)}\mathrm{d}w+\frac{\displaystyle x_{i}f_{X_{i}}(x_{i})}{\displaystyle F_{X_{i}}(x_{i})},$
    $=-\int_{x_{i}}^{\infty}h_{X_{i}}(w)\mathrm{d}w+x_{i}h_{X_{i}}(x_{i}).$
@@
    $x_{i}h_{X_{i}}(x_{i})\leq\int_{x_{i}}^{\infty}h_{X_{i}}(w)\mathrm{d}w,\quad\forall x_{i}\in[0,\infty).$
@@
    $\iff G_{i}(\lambda x_{i})$ $\leq\frac{\displaystyle 1}{\displaystyle\lambda}G_{i}(x_{i}).$
@@
Let $\bm{X}=(X_{1},X_{2},X_{3})$ be an independent random vector (a special case of NSD). Assume that $X_{1}$ and $X_{2}$ are Fréchet distributed with unit scales and shape parameters $\alpha_{1}=\alpha_{2}=\tfrac{1}{2}$, while $X_{3}$ has a piecewise CDF composed of a power-law part followed by a Fréchet CDF with $\theta_{3}=1$ and $\alpha_{3}=\tfrac{1}{2}$. Explicitly,
@@
The corresponding $\phi_{i}$-functions (as defined in Theorem  3.4) are
@@
It is clear that $\phi_{1}$ and $\phi_{2}$ are non-increasing, whereas $\phi_{3}$ fails to be non-increasing on the interval $x_{3}\in\left[\frac{\displaystyle 1}{\displaystyle\sqrt{e}},1\right]$. Figure 2 shows the graphs of these functions.
@@
Figure 2: The marginal $\phi_{i}$ functions.
@@
While the VaR of the sum $S$ can only be computed numerically, the VaRs of the marginals are given explicitly. For $X_{1}$ and $X_{2}$,
@@
${\mathrm{VaR}}_{p}[X_{3}]=\begin{cases}\sqrt{e}\,\sqrt{p},&0<p\leq\frac{\displaystyle 1}{\displaystyle e},\\[5.69054pt] \dfrac{1}{\log^{2}(1/p)},&\frac{\displaystyle 1}{\displaystyle e}<p<1.\end{cases}$
@@
Figure 3 shows that $\bm{X}$ is VaR super-additive. This example illustrates that an NSD vector may have an SD aggregator $\Phi$ without $\Phi$ being globally non-increasing, while still exhibiting VaR super-additivity.
@@
where each $\xi_{i}:[0,\infty)\to[0,\infty)$ is measurable, preserves the property of VaR super-additivity. Specifically, we seek to identify assumptions on the functions $\xi_{i}$ that ensure $\widetilde{\bm{X}}$ remains VaR super-additive whenever the original vector $\bm{X}=(X_{1},\dots,X_{n})$ is already VaR super-additive.
@@
$\leq y_{i}\frac{\displaystyle\xi_{i}^{-1}(x_{i})}{\displaystyle\xi_{i}^{-1}(y_{i})}\log F_{X_{i}}\left(\xi_{i}^{-1}(x_{i})\right),$
@@
We conclude this section by noting that although the NSD and SD properties allow Theorem 3.4 to characterize a broad class of random vectors for which super-additivity is guaranteed, they are not the only indicators of VaR super-additivity. The following example illustrates situations in which VaR is super-additive even when neither NSD nor SD is satisfied.
@@
    We now compare $F_{S}(t)$ with $F_{X_{1}}(t)F_{X_{2}}(t)$. Since $0<\alpha\leq 1$, one checks that
@@
    Hence $F_{S}(t)\geq F_{X_{1}}(t)F_{X_{2}}(t)$ for all $t\geq 0$, meaning that $\bm{X}$ is not NSD.
    From Example 3.9, the functions $\phi_{i}$ are non-increasing for Pareto (II) marginals with $0<\alpha\leq 1$, and therefore $\Phi$ is SD.
@@
    Next, suppose that $\bm{X}$ is a mutually exclusive [Dhaene1999] discrete vector supported on
    $\bigl\{(2^{-k},0):k\geq 1\bigr\}\ \cup\ \bigl\{(0,2^{-k}):k\geq 1\bigr\},$
@@
    $\mathbb{P}\bigl((X_{1},X_{2})=(2^{-k},0)\bigr)=\mathbb{P}\bigl((X_{1},X_{2})=(0,2^{-k})\bigr)=\frac{1}{2^{\,k-1}},\qquad k\geq 1.$
@@
    $F_{X_{1}}(x)=F_{X_{2}}(x)=\begin{cases}0,&x<0,\\[4.0pt] \tfrac{1}{2},&0\leq x<2,\\[4.0pt] 1-2^{-(k+1)},&2^{k}\leq x<2^{k+1},\qquad k\geq 1,\end{cases}$
@@
    so $\bm{X}$ is NSD. This expected as the random vector $\bm{X}$ has a counter-monotonic joint law which belongs to the NSD class.
    The marginal CDFs are discontinuous, and the functions $\phi_{i}(x)=x\log F_{X_{i}}(x)$ are not non-increasing. To see the latter, take $x=2^{k}<y=2^{k+1}$, $k\geq 1$. Then
    $\phi_{i}(y)-\phi_{i}(x)$ $=2^{k+1}\log\Bigl(1-2^{-(k+2)}\Bigr)-2^{k}\log\Bigl(1-2^{-(k+1)}\Bigr)$
@@
    Thus $\phi_{i}$ is strictly increasing along the sequence $\{2^{k}\}$, and consequently $\Phi$ is not SD (take $x=y=2^{k}$ and $s=x+y=2^{k+1}$ then $\Phi(x,y)=\Phi(2^{k},2^{k})<\Phi(2^{k+1},2^{k+1})=\Phi(s,s)$).
@@
    ${\mathrm{VaR}}_{p}[X_{1}]={\mathrm{VaR}}_{p}[X_{2}]$ $=\begin{cases}0,&0<p\leq\tfrac{1}{2},\\[4.0pt] 2^{k},&1-2^{-k}<p\leq 1-2^{-(k+1)},\qquad k\geq 1,\end{cases}$

### v3

@@ 3 VaR super-additivity
    $\phi_{i}^{\prime}(x_{i})$ $=\alpha_{i}\left(\log\!\left(\dfrac{x_{i}}{1+x_{i}}\right)+\dfrac{1}{1+x_{i}}\right)$
    $=\alpha_{i}\left(-\sum_{k=1}^{\infty}\frac{1}{k}\left(\frac{1}{1+x_{i}}\right)^{k}+\dfrac{1}{1+x_{i}}\right)$
    $\leq\alpha_{i}\left(-\frac{1}{1+x_{i}}+\dfrac{1}{1+x_{i}}\right)=0.$
@@
    Log-Power distribution. If
    $F_{X_{i}}(x_{i})=\exp\!\left(-\dfrac{\log(1+x_{i})^{\alpha_{i}}}{x_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i}\in(-\infty,1),$
@@
    $F_{X_{i}}(x_{i})=\dfrac{1}{2}+\dfrac{1}{\pi}\arctan(\alpha_{i}\log(x_{i})),\qquad x_{i}\geq 0,\ \alpha_{i}>0.$
@@
    $\phi_{i}(x_{i})=x_{i}\log\!\left(\dfrac{1}{2}+\dfrac{1}{\pi}\arctan(\alpha_{i}\log(x_{i}))\right).$
@@
    $\phi_{i}^{\prime}(x_{i})=\log F_{X_{i}}(x_{i})+\dfrac{\alpha_{i}}{\pi(1+(\alpha_{i}\log(x_{i}))^{2})\,F_{X_{i}}(x_{i})}.$
@@
    $\phi_{i}^{\prime}(x_{i})=\log F_{X_{i}}(x_{i})+\dfrac{\alpha_{i}\cos^{2}\theta}{\pi F_{X_{i}}(x_{i})}=\log F_{X_{i}}(x_{i})+\dfrac{\alpha_{i}\sin^{2}(\pi F_{X_{i}}(x_{i}))}{\pi F_{X_{i}}(x_{i})}.$
@@
    $\psi_{i}(u)=\log(u)+\dfrac{\alpha_{i}\sin^{2}(\pi u)}{\pi u},\qquad u\in[0,1].$
@@
    $0<\alpha_{i}\leq\inf_{u\in[0,1]}\dfrac{-\pi u\log(u)}{\sin^{2}(\pi u)}\approx 1.0568.$
@@
    $F_{X_{i}}(x_{i})=\frac{1}{\Gamma(\alpha_{i})}\Gamma\!\left(\alpha_{i},\dfrac{\theta_{i}}{x_{i}}\right),\qquad x_{i}\geq 0,\ \alpha_{i},\theta_{i}>0,$
@@
    $\phi_{i}(x_{i})=x_{i}\log\!\left(\frac{1}{\Gamma(\alpha_{i})}\Gamma(\alpha_{i},\frac{\theta_{i}}{x_{i}})\right).$
@@
    $\phi_{i}^{\prime}(x_{i})=\log F_{X_{i}}(x_{i})+\dfrac{\left(\dfrac{\theta_{i}}{x_{i}}\right)^{\alpha_{i}}\exp(-\theta_{i}/x_{i})}{\Gamma(\alpha_{i})\,F_{X_{i}}(x_{i})}.$
@@
    $\psi_{i}(t)=\log\!\left(\frac{1}{\Gamma(\alpha_{i})}\Gamma(\alpha_{i},t)\right)+\dfrac{t^{\alpha_{i}}\exp(-t)}{\Gamma(\alpha_{i},t)}.$
@@
    a Gautschi-type lower bound (Gautschi, , 1959). Thus $\psi_{i}$ is non-increasing with $\psi_{i}(0)=0$, proving $\psi_{i}(t)\leq 0$ on $0,\infty)$.
@@
A shared characteristic across the distributions presented in Example 3.11 is their non-integrability. This property is, in fact, a necessary consequence whenever $\phi_{i}$ is non-increasing.
Proposition 3.12.
If $\phi_{i}(x_{i})=x_{i}\log F_{X_{i}}(x_{i})$ is non-increasing on $0,\infty)$, then $\mathbb{E}[X_{i}]=\infty$.
Proof.
Assume that $\phi_{i}$ is non-increasing and argue by contradiction. Suppose that $\mathbb{E}[X_{i}]<\infty$.
Since $F_{X_{i}}(x_{i})\leq 1$, we have $\log F_{X_{i}}(x)\leq 0$ and hence $\phi_{i}(x_{i})\leq 0$ for all $x_{i}\in[0,\infty)$. Moreover, because $\phi_{i}$ is non-increasing, it starts at some value $\phi_{i}(0)\in(-\infty,0]$ and can only decrease (or remain constant) as $x_{i}$ increases.
As $\mathbb{E}[X_{i}]<\infty$, the standard tail characterization of integrability yields
[…]
We rewrite
$\phi_{i}(x_{i})=x_{i}\log\!\big(1-\overline{F}_{X_{i}}(x_{i})\big).$
Using the expansion $\log(1-u)=-\sum_{k=1}^{\infty}\frac{u^{k}}{k}$ for $u\in(0,1)$, we obtain
$\phi_{i}(x_{i})=-\sum_{k=1}^{\infty}\frac{x_{i}\,\overline{F}_{X_{i}}(x_{i})^{k}}{k}.$
Since $x_{i}\,\overline{F}_{X_{i}}(x_{i})\to 0$, each term in the series converges to zero, and in particular the dominant term $x_{i}\,\overline{F}_{X_{i}}(x_{i})$ vanishes. Consequently,
$\phi_{i}(x_{i})\longrightarrow 0\qquad\text{as }x_{i}\to\infty.$
Thus, $\phi_{i}$ is a non-increasing function taking values in $(-\infty,0]$ and converging to 0 as $x\to\infty$. This can occur only if $\phi_{i}$ increases somewhere, unless $\phi_{i}\equiv 0$, which would imply $F_{X_{i}}\equiv 1$. Both possibilities contradict the assumptions that $\phi_{i}$ is non-increasing and that $X_{i}$ is non-degenerate.
Therefore, $\mathbb{E}[X_{i}]=\infty$. ∎
By Proposition 3.12, a non-increasing $\phi_{i}$ forces $X_{i}$ to have infinite mean and, in particular, unbounded support, underscoring the restrictive nature of this condition. The following proposition characterizes exactly which distributional properties of $X_{i}$ guarantee it, and provides a readily verifiable sufficient condition.
@@
    $x_{i}\,r_{X_{i}}(x_{i})\leq\int_{x_{i}}^{\infty}r_{X_{i}}(w)\,\mathrm{d}w,\quad\forall x_{i}\in[0,\infty).$
    Where $r_{X_{i}}(x_{i})=\dfrac{f_{X_{i}}(x_{i})}{F_{X_{i}}(x_{i})}$ is the reverse hazard rate function (Block et al., , 1998) of the random variable $X_{i}$. In particular, if $x_{i}^{2}\,r_{X_{i}}(x_{i})$ is non-decreasing then $\phi_{i}(x_{i})$ is non-increasing.
@@
    $G_{i}(\lambda x_{i})\leq\dfrac{1}{\lambda}G(x_{i}),\ \quad\forall\lambda\in[1,\infty).$
@@
    $\phi_{i}^{{}^{\prime}}(x_{i})$ $=\log F_{X_{i}}(x_{i})+\dfrac{x_{i}f_{X_{i}}(x_{i})}{F_{X_{i}}(x_{i})},$
    $=-\int_{x_{i}}^{\infty}\dfrac{f_{X_{i}}(w)}{F_{X_{i}}(w)}\,\mathrm{d}w+\dfrac{x_{i}f_{X_{i}}(x_{i})}{F_{X_{i}}(x_{i})},$
    $=-\int_{x_{i}}^{\infty}r_{X_{i}}(w)\,\mathrm{d}w+x_{i}\,r_{X_{i}}(x_{i}).$
@@
    $x_{i}\,r_{X_{i}}(x_{i})\leq\int_{x_{i}}^{\infty}r_{X_{i}}(w)\,\mathrm{d}w,\quad\forall x_{i}\in[0,\infty).$
    Let us now show that the assumption of $x_{i}^{2}\,r_{X_{i}}(x_{i})$ being non-decreasing implies $\phi_{i}^{\prime}(x_{i})\leq 0$, and hence that $\phi_{i}(x_{i})$ is non-increasing.
    Since $x_{i}^{2}\,r_{X_{i}}(x_{i})$ is non-decreasing then for all $w\geq x_{i}$:
    $w^{2}\,r_{X_{i}}(w)\geq x_{i}^{2}\,r_{X_{i}}(x_{i})\quad\iff\quad r_{X_{i}}(w)\geq\dfrac{x_{i}^{2}\,r_{X_{i}}(x_{i})}{w^{2}}.$
    Integrating both sides of the inequality with respect to $w$ yields:
    $\int_{x_{i}}^{\infty}r_{X_{i}}(w)\,\textrm{d}w\geq\int_{x_{i}}^{\infty}\dfrac{x_{i}^{2}\,r_{X_{i}}(x_{i})}{w^{2}}\,\textrm{d}w=x_{i}^{2}\,r_{X_{i}}(x_{i})\,\dfrac{1}{x_{i}}=x_{i}\,r_{X_{i}}(x_{i}),$
    i.e. $\phi_{i}^{\prime}(x_{i})$ is non-positive for all $x_{i}$. Consequently, $\phi_{i}(x_{i})$ is non-increasing, as claimed.
@@
    $\iff G_{i}(\lambda x_{i})$ $\leq\dfrac{1}{\lambda}G_{i}(x_{i}).$
Remark 3.14.
The sufficient condition of non-decreasing $x_{i}^{2}\,r_{X_{i}}(x_{i})$ in Proposition 3.13(i) is particularly convenient in practice, as it allows one to verify the non-increasing property of $\phi_{i}$ directly from the reverse hazard rate, bypassing any direct analysis of $\phi_{i}$ itself. Many distributions beyond those in Example 3.11 satisfy it. For instance, if $X_{i}$ follows a Log-logistic distribution with unit scale and shape $\alpha_{i}>0$, then
$x_{i}^{2}\,r_{X_{i}}(x_{i})=\frac{\alpha_{i}\,x_{i}}{1+x_{i}^{\alpha_{i}}},$
which is non-decreasing if and only if $0<\alpha_{i}\leq 1$.
The condition also admits a natural interpretation via the inverted variable $1/X_{i}$. Since
$h_{1/X_{i}}(x_{i})=\frac{f_{1/X_{i}}(x_{i})}{\overline{F}_{1/X_{i}}(x_{i})}=\frac{(1/x_{i})^{2}\,f_{X_{i}}(1/x_{i})}{F_{X_{i}}(1/x_{i})}=(1/x_{i})^{2}\,r_{X_{i}}(1/x_{i}),$
$h_{1/X_{i}}(x_{i})$ is non-increasing if and only if $x_{i}^{2}\,r_{X_{i}}(x_{i})$ is non-decreasing, meaning the sufficient condition is equivalent to $1/X_{i}$ having a non-increasing hazard rate. This equivalence is not coincidental: the non-increasing hazard rate of $1/X_{i}$ plays a dual role in establishing VaR super-additivity. It governs the right-tail of $X_{i}$ – as implied by Proposition 3.12 – ensuring super-additivity for $p\in(p^{*},1)$, while simultaneously controlling the body of $X_{i}$ (the tail of $1/X_{i}$), so that super-additivity extends to $p\in(0,p^{*})$. This interplay recurs throughout the literature: Chen and Shneer, (2026) exploited the sub-additivity of $\log\overline{F}_{1/X}(x)$, while Arab et al., (2025) worked with that of $F_{1/X}(x)$, both conditions being rooted in the same structural regularity of $1/X_{i}$.
@@
Let $\bm{X}=(X_{1},X_{2},X_{3})$ be an independent random vector (a special case of NSD). Assume that $X_{1}$ and $X_{2}$ are Fréchet distributed with unit scales and shape parameters $\alpha_{1}=\alpha_{2}=\dfrac{1}{2}$, while $X_{3}$ has a piecewise CDF composed of a power-law part followed by a Fréchet CDF with $\theta_{3}=1$ and $\alpha_{3}=\dfrac{1}{2}$. Explicitly,
@@
The corresponding $\phi_{i}$-functions (as defined in Equation (3.2)) are
@@
It is clear that $\phi_{1}$ and $\phi_{2}$ are non-increasing, whereas $\phi_{3}$ fails to be non-increasing on the interval $x_{3}\in\left[\dfrac{1}{\sqrt{e}},1\right]$. Figure 2 shows the graphs of these functions.
@@
Figure 2: Plots of the functions $\phi_{i}$.
@@
While the VaR of the sum $S$ can only be computed numerically, the VaRs of the margins are given explicitly. For $X_{1}$ and $X_{2}$,
@@
${\mathrm{VaR}}_{p}[X_{3}]=\begin{cases}\sqrt{e}\,\sqrt{p},&0<p\leq\dfrac{1}{e},\\[5.69054pt] \dfrac{1}{\log^{2}(1/p)},&\dfrac{1}{e}<p<1.\end{cases}$
@@
The plot shows that $\bm{X}$ is VaR super-additive. This example illustrates that an NSD vector may have an SD aggregator $\Phi$ without $\Phi$ being globally non-increasing, while still exhibiting VaR super-additivity.
@@
where each $\xi_{i}:[0,\infty)\to[0,\infty)$, is measurable and preserves the property of VaR super-additivity. Specifically, we seek to identify assumptions on the functions $\xi_{i}$ that ensure $\widetilde{\bm{X}}$ remains VaR super-additive whenever the original vector $\bm{X}=(X_{1},\dots,X_{n})$ is already VaR super-additive.
@@
$\leq y_{i}\dfrac{\xi_{i}^{-1}(x_{i})}{\xi_{i}^{-1}(y_{i})}\log F_{X_{i}}\left(\xi_{i}^{-1}(x_{i})\right),$
@@
We conclude this section by noting that although the NSD and SD properties allow Theorem 3.5 to characterize a broad class of random vectors, they are not the only indicators of VaR super-additivity. The following example illustrates situations in which $\bm{X}$ is VaR super-additive even when neither NSD nor SD is satisfied.
@@
    We now compare $F_{S}(t)$ with $F_{X_{1}}(t)F_{X_{2}}(t)$. Since $0<\alpha\leq 1$, by the mean value theorem for $t\mapsto t^{\alpha}$ on $\left[\dfrac{1}{1+t},1\right]$, one checks that
@@
    Hence $F_{S}(t)\geq F_{X_{1}}(t)F_{X_{2}}(t)$ for all $t\geq 0$, meaning that $\bm{X}$ is not NSD. In fact, it satisfies the opposite property, namely Positive Simplex Dependence.
    From Example 3.11, the functions $\phi_{i}$ are non-increasing for Pareto (II) margins with $0<\alpha\leq 1$, and therefore $\Phi$ is SD.
@@
    Next, suppose that $\bm{X}$ is a mutually exclusive (Dhaene and Denuit, , 1999) discrete vector supported on
    $\bigl\{(2^{k},0):k\geq 1\bigr\}\ \cup\ \bigl\{(0,2^{k}):k\geq 1\bigr\},$
@@
    $\mathbb{P}\bigl((X_{1},X_{2})=(2^{k},0)\bigr)=\mathbb{P}\bigl((X_{1},X_{2})=(0,2^{k})\bigr)=\frac{1}{2^{\,k+1}},\qquad k\geq 1.$
@@
    $F_{X_{1}}(x)=F_{X_{2}}(x)=\begin{cases}0,&x<0,\\[4.0pt] \dfrac{1}{2},&0\leq x<2,\\[4.0pt] 1-2^{-(k+1)},&2^{k}\leq x<2^{k+1},\qquad k\geq 1,\end{cases}$
@@
    so $\bm{X}$ is NSD. This is expected as the random vector $\bm{X}$ has a counter-monotonic joint law which belongs to the NSD class.
    The marginal CDFs are discontinuous, and the functions $\phi_{i}(x_{i})=x_{i}\log F_{X_{i}}(x_{i})$ are not non-increasing. To see the latter, take $x_{i}=2^{k}<y_{i}=2^{k+1}$, $k\geq 1$. Then
    $\phi_{i}(y_{i})-\phi_{i}(x_{i})$ $=2^{k+1}\log\Bigl(1-2^{-(k+2)}\Bigr)-2^{k}\log\Bigl(1-2^{-(k+1)}\Bigr)$
@@
    Thus $\phi_{i}$ is strictly increasing along the sequence $\{2^{k}\}$, and consequently $\Phi$ is not SD (take $x_{1}=x_{2}=2^{k}$ and $s=x_{1}+x_{2}=2^{k+1}$ then $\Phi(x_{1},x_{2})=\Phi(2^{k},2^{k})<\Phi(2^{k+1},2^{k+1})=\Phi(s,s)$).
@@
    ${\mathrm{VaR}}_{p}[X_{1}]={\mathrm{VaR}}_{p}[X_{2}]$ $=\begin{cases}0,&0<p\leq\dfrac{1}{2},\\[4.0pt] 2^{k},&1-2^{-k}<p\leq 1-2^{-(k+1)},\qquad k\geq 1,\end{cases}$